OpenStudy (anonymous):

Tyler High School has a total of 101 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many girls play sports at this high school? A 23 B 31 C 70 D 80

OpenStudy (mathstudent55):

Let b = number of boys Let g = number of girls total of 101 boys and girls who play sports b + g = 101 the number of boys is 8 more than twice the number of girls b = 2g + 8 Solve the system of equations: b + g = 101 b = 2g + 8

OpenStudy (anonymous):

@mathstudent55 so g=?

OpenStudy (mathstudent55):

Since in hte second equation you see that b = 2g + 8, use the substitution method and replace b in the first equation by 2g + 8: b + g = 101 2g + 8 + g = 101 3g + 8 = 101 3g = 93 g = 31

OpenStudy (anonymous):

@mathstudent55 THank you SO MUCH! :)

OpenStudy (mathstudent55):

yw